# Pivoting Factorization: A Compact Meta Low-Rank Representation of Sparsity for Efficient Inference in Large Language Models

### v2

@@ 1 Introduction
Combining M and PIFA, we present MPIFA—an end-to-end, retraining-free low-rank pruning framework for LLMs. MPIFA significantly outperforms existing low-rank pruning methods, reducing the perplexity gap by 40%-70% on LLaMA2 models (7B, 13B, 70B) and the LLaMA3-8B model  [^1] https://www.llama.com/. Further experiments show that MPIFA consistently achieves superior speedup and memory savings both layerwise and end-to-end compared to semi-structured pruning, while maintaining comparable or even better perplexity. Notably, as shown in Table 3, at $d=32768$, PIFA with 55% density achieves a 2.1$\times$ speedup, whereas various implementations of semi-sparse methods are either slower than the dense linear layer or fail to execute.
@@ 2.1 Connection-wise pruning
    Semi-structured pruning, i.e., N:M sparsity, enforces that in every group of $M$ consecutive elements, $N$ must be zeroed out. This constraint is hardware-friendly and enables optimized acceleration on GPUs like NVIDIA’s Ampere architecture (Mishra et al., 2021). However, semi-structured pruning is constrained by its sparsity pattern, preventing flexible density adjustments and making it inapplicable for acceleration on general GPUs.
@@ 2.2 Low-rank pruning
Low-rank pruning applies matrix decomposition techniques, such as Singular Value Decomposition (SVD), to approximate weight matrices with lower-rank representations, thereby reducing both storage and computational demands. This approach, compatible with any GPU, represents large matrices as products of smaller ones, improving computational efficiency. Recent studies (Hsu et al., 2022; Yuan et al., 2023; Wang et al., 2024; Jaiswal et al., 2024; Saha et al., 2024; Kaushal et al., 2023; Sharma et al., 2023; Qinsi et al., ) highlight the effectiveness of low-rank decomposition in compressing LLMs. However, despite their flexibility, these methods lag behind semi-structured pruning in performance, often leading to a 2$\times$ increase in perplexity (PPL) at the same densities.
@@ 3.1 Motivation: Information Redundancy in Singular Value Decomposition
The number of parameters in the dense weight matrix is $mn$, whereas the total number of parameters in the low-rank matrices is $r(m+n)$. As shown in Figure 1, SVD-based low-rank representations fail to compress the weight matrix when $r$ exceeds half of the matrix dimensions. However, in low-rank decomposition, the orthogonality constraints among singular vectors reduce the effective degrees of freedom. Specifically, for $\vec{U}$ and $\vec{V}^{\mathrm{T}}$, there are $\binom{r}{2}=\frac{r(r-1)}{2}$ unique pairs of singular vectors for each matrix, and each pair imposes one linear constraint due to orthogonality. Together, these constraints reduce the total degrees of freedom by $r(r-1)$.
@@ 3.3 Memory and Computational Cost of PIFA
  • For the PIFA layer (Algorithm 2), the computational cost includes: $\mathbf{Y}_{p}\leftarrow\mathbf{W}_{p}\mathbf{X}$ $(2rnb)$ and $\mathbf{Y}_{np}\leftarrow\mathbf{C}\mathbf{Y}_{p}$ $2br(m-r)$. The total cost is $2rnb+2br(m-r)=2br(m+n-r)$ FLOPs.
@@ 4 Online Error-Accumulation-Minimization Reconstruction
The term $\mathbf{X}\mathbf{X}^{\mathrm{T}}$ can be computed incrementally as $\mathbf{X}\mathbf{X}^{\mathrm{T}}=\sum_{i=1}^{b}\mathbf{x}_{i}\mathbf{x}_{i}^{\mathrm{T}}$, where $\mathbf{x}_{i}$ represents the input of sample $i$. As $\mathbf{X}\mathbf{X}^{\mathrm{T}}\in\mathbb{R}^{n\times n}$, the memory consumption of the online least squares solution remains constant, regardless of the number of calibration samples.
@@ 5.1 Main Result
We further compare MPIFA with 2:4 semi-structured pruning methods: magnitude pruning (Zhu & Gupta, 2017), and two recent state-of-the-art works Wanda (Sun et al., 2024), and RIA (Zhang et al., 2024). According to (Mishra et al., 2021), for 16-bit operands, 2:4 sparse leads to $\sim$44savings in GPU memory. Therefore, we compare 2:4 sparse method with MPIFA at 0.55 density to ensure that all methods achieve the same memory reduction (see Table 3 for memory comparison).
@@ 5.2 Inference Speedup and Memory Reduction
Figure 4 and Table 3 compare the speedup and memory usage of the PIFA layer with 2:4 semi-sparse linear layers. The semi-sparse layers are implemented using cuSPARSELt [^2] https://docs.nvidia.com/cuda/cusparselt/ or CUTLASS [^3] https://github.com/NVIDIA/cutlass library, with the reported speedup representing the higher value between the two implementations. The results span various dimensions on A6000 and A100 GPUs. PIFA demonstrates consistently superior efficiency, achieving the highest speedup and lowest memory usage in all configurations except $d=4096$. Notably, PIFA’s acceleration increases with matrix dimensions, reflecting its scalability and computational effectiveness. As shown in Table 3, at $d=32768$, PIFA with 55% density achieves a 2.1$\times$ speedup, while 2:4 (CUTLASS) is slower than the dense linear layer, and 2:4 (cuSPARSELt) raises an error.
@@ Acknowledgements
Figure 7: Efficiency of PIFA layer under various ranks, with sequence length = 2048, batch size = 32, and dimension = 8192 on FP32 and FP16 on A6000 GPU. At $50\%$ density, PIFA achieves $47.6\%$ memory savings and 1.95×1.95$\times$1.95 × speedup on FP32. These results guarantee that the overhead of both time and memory of the PIFA layer is quite low.
@@
Table 3: Efficiency of PIFA layer and semi-sparse layer across different dimensions, compared to dense linear at same dimension on same GPU, with sequence length of 2048 and batch size of 32, using FP16 (FP32 is not supported by 2:4 sparsity in torch.sparse). The highest speedup and lowest memory are indicated in bold. PIFA shows increasing speedup as the dimension grows. For matrix multiplication with weight matrix shape 32768$\times$32768, cuSPARSELt raises CUDA error.
@@
Speedup A6000 2:4 (cuSPARSELt) Error 0.94$\times$ 0.97$\times$ 1.09$\times$
2:4 (CUTLASS) 0.79$\times$ 0.92$\times$ 1.15$\times$ 1.18$\times$
PIFA 55% 2.10$\times$ 1.88$\times$ 1.70$\times$ 1.43$\times$
A100 2:4 (cuSPARSELt) Error 1.19$\times$ 1.31$\times$ 1.68$\times$
2:4 (CUTLASS) 1.19$\times$ 1.12$\times$ 1.09$\times$ 1.52$\times$
PIFA 55% 1.98$\times$ 1.70$\times$ 1.54$\times$ 1.37$\times$
@@ A.1 Direct Optimization
$-2\mathbf{A}+2\mathbf{B}\mathbf{V}\mathbf{C}=0\implies\mathbf{B}\mathbf{V}\mathbf{C}=\mathbf{A}$
@@ B.2 MPIFANS
[…]
@@
Module Density=Type Density×Layer DensityGlobal $\text{Module Density}=\frac{\text{Type Density}\times\text{Layer Density}}{\text{Global Density}}.$

### v3

@@ 1 Introduction
[…]
@@ 2.1 Connection-wise pruning
    Semi-structured pruning, i.e., N:M sparsity, enforces that in every group of $M$ consecutive elements, NN must be zeroed out. This constraint is hardware-friendly and enables optimized acceleration on GPUs like NVIDIA’s Ampere architecture (Mishra et al., 2021). However, semi-structured pruning is constrained by its sparsity pattern, preventing flexible density adjustments and making it inapplicable for acceleration on general GPUs.
@@ 2.2 Low-rank pruning
[…]
@@ 3.1 Motivation: Information Redundancy in Singular Value Decomposition
The number of parameters in the dense weight matrix is mnmn, whereas the total number of parameters in the low-rank matrices is $r(m+n)$. As shown in Figure 1, SVD-based low-rank representations fail to compress the weight matrix when $r$ exceeds half of the matrix dimensions. However, in low-rank decomposition, the orthogonality constraints among singular vectors reduce the effective degrees of freedom. Specifically, for $\vec{U}$ and $\vec{V}^{\mathrm{T}}$, there are $\binom{r}{2}=\frac{r(r-1)}{2}$ unique pairs of singular vectors for each matrix, and each pair imposes one linear constraint due to orthogonality. Together, these constraints reduce the total degrees of freedom by r(r1)r(r-1).
@@ 3.3 Memory and Computational Cost of PIFA
  • For the PIFA layer (Algorithm 2), the computational cost includes: $\mathbf{Y}_{p}\leftarrow\mathbf{W}_{p}\mathbf{X}$ $(2rnb)$ and $\mathbf{Y}_{np}\leftarrow\mathbf{C}\mathbf{Y}_{p}$ 2br(mr)2br(m-r). The total cost is 2rnb+2br(mr)=2br(m+nr)2rnb+2br(m-r)=2br(m+n-r) FLOPs.
@@ 4 Online Error-Accumulation-Minimization Reconstruction
The term $\mathbf{X}\mathbf{X}^{\mathrm{T}}$ can be computed incrementally as $\mathbf{X}\mathbf{X}^{\mathrm{T}}=\sum_{i=1}^{b}\mathbf{x}_{i}\mathbf{x}_{i}^{\mathrm{T}}$, where $\mathbf{x}_{i}$ represents the input of sample ii. As $\mathbf{X}\mathbf{X}^{\mathrm{T}}\in\mathbb{R}^{n\times n}$, the memory consumption of the online least squares solution remains constant, regardless of the number of calibration samples.
@@ 5.1 Main Result
We further compare MPIFA with 2:4 semi-structured pruning methods: magnitude pruning (Zhu & Gupta, 2017), and two recent state-of-the-art works Wanda (Sun et al., 2024), and RIA (Zhang et al., 2024). According to (Mishra et al., 2021), for 16-bit operands, 2:4 sparse leads to $\sim$44% savings in GPU memory. Therefore, we compare 2:4 sparse method with MPIFA at 0.55 density to ensure that all methods achieve the same memory reduction (see Table 3 for memory comparison).
@@ 5.2 Inference Speedup and Memory Reduction
[…]
@@ Acknowledgements
Figure 7: Efficiency of PIFA layer under various ranks, with sequence length = 2048, batch size = 32, and dimension = 8192 on FP32 and FP16 on A6000 GPU. At $50\%$ density, PIFA achieves $47.6\%$ memory savings and $1.95\times$ speedup on FP32. These results guarantee that the overhead of both time and memory of the PIFA layer is quite low.
@@
Table 3: Efficiency of PIFA layer and semi-sparse layer across different dimensions, compared to dense linear at same dimension on same GPU, with sequence length of 2048 and batch size of 32, using FP16 (FP32 is not supported by 2:4 sparsity in torch.sparse). The highest speedup and lowest memory are indicated in bold. PIFA shows increasing speedup as the dimension grows. For matrix multiplication with weight matrix shape $\times32768$, cuSPARSELt raises CUDA error.
@@
[…]
@@ A.1 Direct Optimization
$2\mathbf{A}+2\mathbf{B}\mathbf{V}\mathbf{C}=0\implies\mathbf{B}\mathbf{V}\mathbf{C}=\mathbf{A}$
@@ B.2 MPIFANS
Type Density introduces non-uniform sparsity between attention and MLP modules. Based on insights from prior literature (Yuan et al., 2023), MLP modules exhibit higher sensitivity to pruning compared to attention modules. To account for this, we search for the density of attention modules within $\{\text{Global Density},\text{Global Density}-0.1\}$, optimizing for performance. The density of MLP modules is then calculated to ensure that the model’s global density remains unchanged.
@@
$\text{Module Density}=\frac{\text{Type Density}\times\text{Layer Density}}{\text{Global Density}}.$